# Finding Dori [Uncaptioned image]: Memorization in Text-to-Image Diffusion Models Is Less Local Than Assumed

### v1

@@ 4.2 Images are Not Memorized in a Subset of Weights
The activations are collected from the layers where NeMo and Wanda operate: the cross-attention modules’ value layer for NeMo and their output fully-connected layers for Wanda. We compute activations for seven cross-attention modules, indexed from 1 to 7, spanning the three Down blocks (indices 1 to 6, each block has two modules) and the Mid block of the U-Net [34], following the setup of NeMo.
[…]
While high discrepancy scores suggest that different weights may contribute to data replication, we investigate this further by analyzing the consistency of pruning-based mitigation methods across different adversarial embeddings for the same image. Since NeMo and Wanda rely on activations to select weights for pruning, we expect these sets to vary across different adversarial embeddings. To validate this expectation, we define the weight agreement between two adversarial embeddings as the overlap of identified weights from NeMo or Wanda, normalized by the total number of identified weights, i.e., the intersection over union. The final agreement metric is computed as the average pairwise weight agreement across a set of adversarial embeddings. The precise formulation of this metric is provided in Eq. 5 in Sec. G.2. If no weights are identified for a pair in a given layer, we set the agreement to 1. As with the discrepancy metric, we report the mean agreement per layer.
[…]
Low weight agreement, high activation discrepancy, and the abundance of adversarial embeddings capable of replicating any memorized image suggest that memorization in DMs is a more complex phenomenon than previously assumed, and that local interventions are insufficient to address it. However, we refrain from claiming that our results definitively disprove the existence of locality. Rather, we emphasize the inherent difficulty in reliably identifying the specific weights responsible for memorization, assuming such weights exist. Since current methods rely on activation patterns to detect memorization-related weights, they fail when adversarial embeddings induce divergent activations, leading to inconsistent weight selection. Moreover, in an attempt to remove memorized content, these methods often need to prune so many weights that the model becomes virtually unusable. This behavior suggests that effectively eliminating memorized content from a DM may require interventions that affect all weights, rather than targeting only a small subset through pruning.
Figure 3: Locality through the lens of activations and memorization weights. Although adversarial embeddings trigger the same image, their activations (left) exhibit high discrepancy—comparable to that of embeddings causing replication of different images. Since memorization weight identification relies on activations, this large discrepancy results in low weight agreement (right), which undermines the idea that weights responsible for replicating a memorized image can be pinpointed and pruned.
5 Robust Mitigation via Adversarial Fine-Tuning
To address the limitations of pruning-based approaches, we propose adversarial fine-tuning as a robust method for permanently removing memorized content from DMs.
5.1 Adversarial Fine-Tuning
Building on our earlier findings that pruning-based mitigation methods conceal rather than eliminate memorization, we propose a fine-tuning-based approach designed to permanently remove memorized content from the model. Our method takes inspiration from the concept of adversarial training of image classifiers [40, 10, 26], which iteratively crafts adversarial examples and updates the classifier to resist them. Our fine-tuning is inspired by the adversarial perspective in Sec. 3, in which an adversary iteratively tries to trigger undesired data replication of memorized content.
Before fine-tuning, we first generate a set of images conditioned on the memorized prompt while applying a mitigation method that suppresses data replication to collect a set of surrogate samples. We refer to individual generated surrogate images as $\widetilde{{\bm{x}}}_{0}$, with their corresponding noisy versions at diffusion time step $t=T,\ldots,1$ as $\widetilde{{\bm{x}}}_{t}$. Since the generation of these surrogate images is independent of the fine-tuning process, images can be generated beforehand.
During each fine-tuning step, we first sample a memorized image and search for a corresponding trigger embedding ${\bm{y}}_{\text{adv}}$, following our procedure described in Sec. 3.1. We switch between initializing the embedding ${\bm{y}}_{\text{adv}}^{\smash{(0)}}$ with the memorized prompt’s embedding and drawing from a random Gaussian distribution to increase robustness to adversarial embeddings crafted from different initializations. Based on these trigger embeddings, we fine-tune the noise predictor $\bm{\epsilon}_{\bm{\theta}}$ to disrupt trajectories that would otherwise lead to data replication by steering it toward surrogate trajectories. For each batch, we sample multiple time steps $t$ and use the corresponding surrogate noisy images $\widetilde{{\bm{x}}}_{t}$ to update the model. We fine-tune the model using the standard diffusion training objective, conditioned on the adversarial embedding:
$\mathcal{L}_{Adv}(\widetilde{{\bm{x}}}_{0},\bm{\epsilon},{\bm{y}}_{adv},t,\bm{\theta})=\|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}\left(\widetilde{{\bm{x}}}_{t},t,{\bm{y}}_{adv}\right)\|_{2}^{2}.$ (3)
Fine-tuning exclusively on adversarial embeddings and generated images risks degrading the model’s general utility. To preserve performance on non-memorized data, we train the model on additional non-memorized image-captions pairs using the standard diffusion loss $\mathcal{L}_{\text{DM}}$ defined in Eq. 1. In practice, we perform a single forward pass per update step by concatenating the embeddings and corresponding images used for both the adversarial and standard loss components to speed up the training. An algorithmic description of our adversarial fine-tuning method is provided in Appx. F.
5.2 Experimental Results of Adversarial Fine-tuning
As before, we rely on Stable Diffusion v1.4. In this section, we focus on removing VM samples from the model, as the impact on this type of memorization can be clearly quantified using SSCD-based metrics. We use a subset of the LAION-Aesthetics [36] dataset as non-memorized samples to compute the second loss. Throughout our experiments, we use NeMo to generate 25 surrogate samples per memorized prompt. However, we emphasize that any effective mitigation method—whether through pruning model components or modifying inputs at inference time—is suitable. As the goal is to preserve semantic alignment while removing memorization, the surrogate samples can also be generated from a separate diffusion model, which does not replicate the target example.
The DM’s U-Net is fine-tuned for up to 50 epochs using the Adam optimizer with a learning rate of 1e-5. For simplicity, we use the same batch size of 4 for both surrogate and non-memorized samples, keeping a 1:1 ratio. After crafting an adversarial embedding, we perform three consecutive update steps with the same embedding but different image batches to reduce computational cost. Adversarial embeddings are generated following the hyperparameters described in Sec. 3.2.
We find that our adversarial fine-tuning procedure quickly removes memorized content. Tab. 1 (bottom row) presents the evaluation results after fine-tuning for five epochs. The results show that adversarial embeddings can no longer trigger data replication, indicating a permanent mitigation compared to pruning-based methods. At the same time, the model’s utility is preserved: the FID score improves from 14.44 to 13.61 after fine-tuning, suggesting no harm to the image quality.
We conduct an extensive analysis of how the model’s robustness evolves during fine-tuning and find that it remains robust after five epochs. While longer training continues to improve evaluation metrics, in practice, five epochs are sufficient to erase memorized content. This result also holds when increasing the number of optimization steps during the search for adversarial embeddings at evaluation time. Additional sensitivity analysis results are reported in Appx. F. Overall, our novel adversarial fine-tuning method is the first to achieve permanent and robust mitigation of undesired memorization in DMs after their initial training.
6 Conclusions
We demonstrate that memorization in text-to-image DMs is not strictly a local phenomenon. While pruning-based methods such as NeMo and Wanda can suppress the generation of a memorized training image when prompted with its original caption, they do not remove the underlying memorization from the model. As a result, the same image can still be regenerated by optimizing the prompt embedding. In effect, pruning alters the generation trajectory for the original prompt, but adversarial embeddings can steer the denoising process along alternative paths that nonetheless reproduce the memorized content. Building on this insight, we propose a novel fine-tuning strategy that leverages adversarial embeddings to fully erase memorized samples from the model. Our work represents a concrete step toward the responsible deployment of generative models by enabling a deeper understanding of memorization and offering a reliable mechanism for its removal.
Acknowledgments and Disclosure of Funding
This research was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project number 550224287. This work has been financially supported by the German Research Center for Artificial Intelligence (DFKI) project “SAINT”. Responsibility for the content of this publication lies with the authors.
@@ Appendix B Hard- and Software Details
We conducted all experiments on NVIDIA DGX systems running NVIDIA DGX Server Version 5.2.0 and Ubuntu 20.04.6 LTS. The machines are equipped with 2 TB of RAM and feature NVIDIA A100-SXM4-40GB GPUs. The respective CPUs are AMD EPYC 7742 64-core. Our experiments utilized CUDA 12.2, Python 3.10.13, and PyTorch 2.2.0 with Torchvision 0.17.0 [28]. Notably, all experiments are conducted on single GPUs.
@@ Appendix C Model and Dataset Details
Our experiments primarily use Stable Diffusion v1-4 [33], which is publicly available at https://huggingface.co/CompVis/stable-diffusion-v1-4. Comprehensive information about the data, training parameters, limitations, and environmental impact can be found at that URL. The model is released under the CreativeML OpenRAIL M license.
@@
The memorized prompts analyzed in our study originate from the LAION2B-en [36] dataset, which was used to train the DM. We use a set of memorized prompts provided by Wen et al. [45] [^1] Available at https://github.com/YuxinWenRick/diffusion_memorization., who identified them using the extraction tool developed by Webster [43]. The LAION dataset is licensed under the Creative Commons CC-BY 4.0. As the images in the dataset may be subject to copyright, we do not include them in our codebase; instead, we provide URLs that allow users to retrieve the images directly from their original sources. For performing our fine-tuning-based mitigation method, we furthermore downloaded 100k images from the LAION aesthetics dataset, a subset of LAION5B.
Appendix D Additional Details and Experiments on Adversarial Embedding Optimization
[…]
@@ Appendix D Additional Details and Experiments on Adversarial Embedding Optimization
      Diffusion model $\bm{\epsilon}_{\bm{\theta}}$triangleright optionally after pruning-based mitigation applied
@@
      Number of optimization steps $N$
@@
${\bm{y}}_{\text{adv}}^{(0)}\leftarrow\text{encode\_text}({\bm{p}}_{\text{mem}})$triangleright alternatively, initialize ${\bm{y}}_{\text{adv}}^{(0)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@
     $t\sim\text{Uniform}(\{1,\ldots,T\})$triangleright sample discrete timestep from noise schedule
     $\tilde{{\bm{x}}}_{t}\leftarrow\text{add\_noise}({\bm{x}}_{\text{mem}},\bm{\epsilon},t)$triangleright adding noise using the training noise scheduler
@@
     ${\bm{y}}_{\text{adv}}^{(i)}\leftarrow{\bm{y}}_{\text{adv}}^{(i-1)}-\eta\cdot\nabla_{{\bm{y}}_{\text{adv}}^{(i-1)}}\|\bm{\epsilon}-\hat{\bm{\epsilon}}\|_{2}^{2}$triangleright update adv. embedding with gradient descent
@@
D.1 Can We Make a DM to Output Any Image With Adversarial Embeddings?
To assess whether Dori’s ability to replicate memorized images is truly due to memorization, we also test whether adversarial text embeddings can be used to generate an arbitrary (non-memorized) image, as described in Sec. 3.1. Intuitively, we expect that we can force an 800M parameter model to produce a specific output vector (latent representation of an image) of size 16,384, given we perform an unconstrained gradient-based optimization of an input vector (text embedding) of size 59,136. In effect, if the model can be forced to produce arbitrary, non-memorized images using these embeddings, it would suggest that Dori is not exploiting memorization, but rather steering the model toward designated outputs—regardless of whether the content was memorized.
To generate non-memorized images with Dori, we sample 100 images from the COCO2014 training set and run the optimization for 1000 steps for each sampled image. Using the resulting adversarial embeddings, we generate 5 images per embedding with SD-v1.4 and compute the SSCD scores between the generated and original images. The SSCD scores for all examples exceed the memorization threshold of 0.7, and qualitatively, Fig. 4 shows that the images are replicated almost perfectly.
While this initially might seem as if Dori is not only replicating memorized samples, we demonstrate in Sec. D.2 that there is, in fact, a difference between triggering generation of memorized versus non-memorized content.

### v2

@@ 4.4 Abandoning the Locality Assumption Yields Robust Mitigation Method
ESD + [Uncaptioned image] $0.90\pm 0.04$ 0.98
Concept Ablation + [Uncaptioned image] $0.91\pm 0.04$ 0.97
SISS + [Uncaptioned image] $0.60\pm 0.22$ 0.39
Our Mitigation + [Uncaptioned image] $\bm{0.36}\pm 0.14$ $\bm{0.02}$
Baseline. We compare our method with the state-of-the-art fine-tuning approach for data point unlearning in DMs, namely SISS (Alberti et al., ), which aims to remove a specific (not necessarily memorized) image from the pre-trained DM. We remove a single image from the U-Net by performing 35 update steps. For each image, we start from the original DM, following the default setting for SISS. We describe SISS and its setup in Section˜E.2. While our results in Tab.˜2 show that it successfully drops $\text{SSCD}_{\text{Orig}}$ below the memorization threshold of $0.7$, we note that SISS fails to remove 39% of memorized samples, as indicated by the memorization rate (MR). This highlights that the method is still limited for the reliable mitigation of memorized images when faced with adversarial embeddings.
Applying Concept Unlearning. Finally, for completeness, we also evaluate ESD (Gandikota et al., 2023) and Concept Ablation (Kumari et al., 2023), state-of-the-art concept unlearning methods for DMs. As discussed in Sec.˜2, concept unlearning successfully targets the removal of broad, high-level concepts from generative models, such as style or entire object categories, but it is not designed to remove specific, individual memorized data points, which is the goal of memorization mitigation. Nevertheless, as the closest available baselines from concept unlearning, we include these approaches in our evaluation. As shown in Tab.˜2 (top two rows), and Tab.˜3 (see Section˜G.2), both methods fail to reliably prevent the replication of memorized data when confronted with adversarial trigger embeddings. We hypothesize this is because these methods rely on a single prompt, or its augmented versions, as a replication trigger, which is insufficient for thorough memorization mitigation, as we already demonstrated in Section˜4.1. We extend the discussion and experimental setup in Appx.˜G.
Overall, our results highlight that existing concept and data unlearning methods are ill-suited for mitigating memorization. Although concept unlearning methods can suppress broad categories, they fail to reliably eliminate memorized content. Similarly, the current state-of-the-art data unlearning method (SISS), which claims to be effective at removing memorized images, also remains vulnerable to adversarial embeddings. Moreover, while existing mitigation methods are effective at preventing data replication when triggered from the prompt space, fully removing memorization requires novel approaches. Such methods must abandon the locality assumption in input and weight space, and operate globally, for instance, through adversarial fine-tuning with many triggers, as in our approach.
5 Discussion and Conclusions
Our findings reveal that memorization in text-to-image DMs is not inherently a local phenomenon. While pruning-based methods such as NeMo and Wanda can suppress the generation of memorized training images when prompted with the original captions, they do not remove the image memorization from the model. In particular, we show that the memorized images can still be reliably regenerated using diverse adversarial embeddings. While our results strongly indicate the limitations of targeting only local weights or activation patterns, we refrain from claiming that any form of locality can be definitively ruled out. Rather, our work highlights the significant challenge in reliably identifying which specific weights drive memorization, especially since current methods based on activation patterns fail in the presence of adversarial triggers that induce divergent activations. We show that effective memorization removal requires global model interventions, as realized by our adversarial fine-tuning approach, which robustly eliminates memorized images while preserving the model’s generative performance. Overall, our insights underscore the need for novel global, model-wide memorization mitigation strategies to support the responsible deployment of generative models.
Reproducibility Statement
For reproducibility, we describe all settings, models, datasets, and hyperparameters used during our experiments. Our Appendix further provides detailed descriptions for all steps of our method. All datasets, models, and methods used in this paper are publicly available.
@@ Appendix A Limitations
We focus our research only on one model, Stable Diffusion v1.4, since only for this DM a set of memorized images is currently known. We acknowledge that such a narrow scope limits the generalizability of our results. However, we would like to point out that other works successfully advanced the understanding of memorization by analyzing Stable Diffusion v1.4.
Misuse of Dori is possible. An adversary with full access to the model and a (suspected) memorized image can perform optimization to trigger replication of the image. We would like to argue that such a scenario is highly unlikely, since the access assumptions for the adversary are very strong, and hard to meet in practice. We designed Dori as an assessment tool to evaluate whether (previously) memorized content remains present in the DM.
@@ Appendix B Hard- and Software Details
We conducted all experiments on NVIDIA DGX systems running NVIDIA DGX Server Version 5.2.0 and Ubuntu 20.04.6 LTS. The machines are equipped with 2 TB of RAM and feature NVIDIA A100-SXM4-40GB GPUs. The respective CPUs are AMD EPYC 7742 64-core. Our experiments utilized CUDA 12.2, Python 3.10.13, and PyTorch 2.2.0 with Torchvision 0.17.0 Paszke et al. (2019). Notably, all experiments are conducted on single GPUs.
@@ Appendix C Model and Dataset Details
Our experiments primarily use Stable Diffusion v1-4 (Rombach et al., 2022), which is publicly available at https://huggingface.co/CompVis/stable-diffusion-v1-4. Comprehensive information about the data, training parameters, limitations, and environmental impact can be found at that URL. The model is released under the CreativeML OpenRAIL M license.
@@
The memorized prompts analyzed in our study originate from the LAION2B-en (Schuhmann et al., 2022) dataset, which was used to train the DM. We use a set of memorized prompts provided by Wen et al. (2024) [^1] Available at https://github.com/YuxinWenRick/diffusion_memorization., who identified them using the extraction tool developed by Webster (2023). The LAION dataset is licensed under the Creative Commons CC-BY 4.0. As the images in the dataset may be subject to copyright, we do not include them in our codebase; instead, we provide URLs that allow users to retrieve the images directly from their original sources. For performing our fine-tuning-based mitigation method, we furthermore downloaded 100k images from the LAION aesthetics dataset, a subset of LAION5B.
Appendix D Declaration on Large Language Models’ (LLMs’) Usage
LLMs were only used to assist in writing (grammar check, phrasing), and to implement boilerplate, repetitive code for data processing and visualizations. No novelty of our work came from an LLM assistant. All new ideas and methods described in the paper are developed by the authors without the assistance of any AI system.
Appendix E Extended Background
E.1 Text-to-Image Diffusion Models
We present the technical details of DM training: During training, a time step $t\sim\mathcal{U}(1,T)$ and a noise vector $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ are randomly sampled to create a noisy image ${\bm{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\bm{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ based on the training image ${\bm{x}}_{0}$. The amount of noise added is controlled by a noise scheduler $\bar{\alpha}_{t}$, for which there are multiple choices (Song & Ermon, 2020; Nichol & Dhariwal, 2021; Kingma et al., 2021; Karras et al., 2022). The training objective for the noise predictor $\bm{\epsilon}_{\bm{\theta}}$ is then to predict the noise $\bm{\epsilon}$ that has been added:
Training and generating samples with DMs can be computationally expensive. The latent DM framework (Rombach et al., 2022) reduces this burden by operating in a lower-dimensional latent space instead of the pixel space. This latent space is learned by a separately trained variational autoencoder (Kingma et al., 2013; Van Den Oord et al., 2017) that encodes images into compact representations and decodes generated latents back to the image space.
E.2 Subtracted Importance Sampled Scores (SISS)
Alberti et al. propose a novel data unlearning method, which aims to remove arbitrary images from the DM, while retaining overall generative capabilities. They perform a full fine-tuning on the U-Net, with minimization objective $\mathcal{L}_{s,\lambda}(\theta)$, defined as:
$\begin{split}\mathbb{E}_{p_{X}(x)}\mathbb{E}_{p_{A}(a)}\mathbb{E}_{q_{\lambda}(m_{t}|x,a)}\Biggl[&\frac{n}{n-k}\frac{q(m_{t}|x)}{(1-\lambda)q(m_{t}|x)+\lambda q(m_{t}|a)}\left\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}\\ &-(1+s)\frac{k}{n-k}\frac{q(m_{t}|a)}{(1-\lambda)q(m_{t}|x)+\lambda q(m_{t}|a)}\left\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}\Biggr]\end{split}$ (4)
Here $X$ denotes the subset of the DM’s training data of size $n$, $A\subset X$ a set of images to unlearn of size $k$, $q_{\lambda}(m_{t}|x,a):=(1-\lambda)q(m_{t}|x)+\lambda q(m_{t}|a)$ is a mixture distribution—a weighted average of data densities $q(m_{t}|x)$ and $q(m_{t}|a)$ parameterized by $\lambda\in[0;1]$, and $\gamma_{t}$ and $\sigma_{t}$ are the DDPM (Ho et al., 2020) forward process parameters at timestep $t$. $\mathcal{L}_{s,\lambda}(\theta)$ provides a middle ground between naive deletion, i.e., fine-tuning only on $X\setminus A$, and NegGrad (Golatkar et al., 2020), which performs gradient ascent on $A$, but is known for its instability. Parameter $\lambda$, with the default value of $0.5$ in their work regulates how much SISS resembles the former and the latter removal methods. To increase the strength of the method, a superfactor hyperparameter $s>0$ is introduced.
For SISS to work we need an access to the subset of the training data $X$, and images to remove $A$.
Applying SISS to Memorization Mitigation is straightforward. We use the default setting specified in the paper for Stable Diffusion v-1.4, i.e., we apply SISS on a single memorized image at a time. Authors provide sets $A$ and $X$ for a small subset of 33 VM images they attempt to remove from the DM in their work. We extend them to the remaining VM samples for a fair comparison with other methods. Specifically, following their approach we add random tokens to the memorized prompt and generate 128 images from the original DM, where $A$ would consist of replicas of the memorized image, while $X$ would contain the remaining (non-memorized) images. We note that the method employed by the authors (originally proposed by Wen et al. (2024)) is unreliable in creating prompts that can generate both novel and memorized images, and for 9 of the memorized samples we fail to craft such (partially memorized) prompts even after varying the number of added tokens. Effectively, 9 out of 112 VM images remain in the model, since we are not able to provide the SISS method with the necessary $X$ and $A$ sets. Notably, the $X$ set is akin to set of surrogate samples used in our mitigation method, which we obtain reliably with a weak, pruning-based removal method (NeMo).
Once we have access to $X$ and $A$, we run SISS with the default hyperparameters for 35 update steps, with learning rate of $10^{-5}$, batch size of 1 and gradient accumulation of 16. For each image we start from the original U-Net, following the setup in the original work. We note that such a setup has a limited applicability in the real-world use-cases, as some DMs (specifically: SD-v1.4) may memorize more than one image.
Appendix F Additional Details and Experiments on Adversarial Embedding Optimization
In the following, we elaborate on the design of the adversarial optimization Eq. (2) used to obtain ${\bm{y}}_{adv}$ to trigger generation of ${\bm{x}}_{\mathit{mem}}$. First, we provide the algorithm in Alg.˜1. Then we showcase that naive unconstrained optimization would yield False Positives (${\bm{y}}_{adv}$ that are capable of forcing the DM to generate arbitrary images), see Section˜F.1. Motivated by this finding, we experiment with the varying strength of the optimization, and arrive at the final constraint of $50$ optimization steps, which allows us to successfully craft ${\bm{y}}_{adv}$ if the optimization target (image) is memorized, and fail to provide ${\bm{y}}_{adv}$ for all other (non-memorized) targets. We evaluate constraining schemes that work in the embedding space in Section˜F.3, and show that they are unsuccessful at preventing False Positives.
@@ Appendix F Additional Details and Experiments on Adversarial Embedding Optimization
  DM $\bm{\epsilon}_{\bm{\theta}}$ $\triangleright$ optionally after pruning-based mitigation applied
@@
   Number of optimization steps NN
@@
${\bm{y}}_{\text{adv}}^{(0)}\leftarrow\text{encode\_text}({\bm{p}}_{\text{mem}})$ $\triangleright$ alternatively, initialize ${\bm{y}}_{\text{adv}}^{(0)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
@@
  $t\sim\text{Uniform}(\{1,\ldots,T\})$ $\triangleright$ sample discrete timestep from noise schedule
  $\tilde{{\bm{x}}}_{t}\leftarrow\text{add\_noise}({\bm{x}}_{\text{mem}},\bm{\epsilon},t)$ $\triangleright$ adding noise using the training noise scheduler
@@
  ${\bm{y}}_{\text{adv}}^{(i)}\leftarrow{\bm{y}}_{\text{adv}}^{(i-1)}-\eta\cdot\nabla_{{\bm{y}}_{\text{adv}}^{(i-1)}}\|\bm{\epsilon}-\hat{\bm{\epsilon}}\|_{2}^{2}$ $\triangleright$ update adv. embedding with gradient descent
@@
F.1 Can We Make a DM to Output Any Image With Adversarial Embeddings?
To assess whether Dori’s ability to replicate memorized images is truly due to memorization, we also test whether adversarial text embeddings can be used to generate an arbitrary (non-memorized) image, as described in Sec.˜3.1. Intuitively, we expect that we can force an 800M parameter model to produce a specific output vector (latent representation of an image) of size 16,384, given we perform an unconstrained gradient-based optimization of an input vector (text embedding) of size 59,136. In effect, if the model can be forced to produce arbitrary, non-memorized images using these embeddings, it would suggest that Dori is not exploiting memorization, but rather steering the model toward designated outputs—regardless of whether the content was memorized.
To generate non-memorized images with Dori, we sample 100 images from the COCO2014 training set and run the optimization for 1000 steps for each sampled image. Using the resulting adversarial embeddings, we generate 5 images per embedding with SD-v1.4 and compute the SSCD scores between the generated and original images. The SSCD scores for all examples exceed the memorization threshold of 0.7, and qualitatively, Fig.˜5 shows that the images are replicated almost perfectly.
While this initially might seem as if Dori is not only replicating memorized samples, we demonstrate in Section˜F.2 that there is, in fact, a difference between triggering generation of memorized versus non-memorized content.